# PoliPrompt: A High-Performance Cost-Effective LLM-Based Text Classification Framework for Political Science

### v1

@@ 3.1 Pre-process
With natural language inputs, LLMs can convert them into numerical embedding vectors of uniform size. We apply this process to all unlabeled texts in the dataset using embedding models such as “text-embedding-3-small” from OpenAI. Based on relevant literature (Steck_2024), we use cosine distance to measure the similarity between embedding pairs, providing a quantitative basis for comparing text semantics.
@@
Uniform Manifold Approximation and Projection (UMap) is a powerful dimensionality reduction technique that preserves data relationships in lower-dimensional space. We employ UMAP (mcinnes2020umapuniformmanifoldapproximation) with cosine distance for feature reduction of the embeddings, which offers two key advantages: it allows us to use simpler Euclidean distance calculations on the reduced embeddings in subsequent stages, and it significantly decreases the feature set size, thereby accelerating computational processes.
Exemplar selection.
To create a diverse and representative pool of examples, we employ an exemplar selection method (Bien_2011) based on the manifold structure of the reduced embeddings. While various approaches exist for exemplar or prototype selection, such as set cover algorithms and density-based sampling, we opt for the k-means selector due to its simplicity and effectiveness. This method involves performing k-means clustering on the embeddings and then designating the text whose embedding is nearest to each cluster center as an exemplar. This approach ensures that every distinct group within the data is represented by one example, allowing for easy control over the number of exemplars selected while maintaining a comprehensive coverage of the embedding space. We keep $M$ such texts as an exemplar pool and engage a human expert to label them. Typically, M is less than 100, requiring minimal human effort.
3.2 In-process
In this stage, we use the exemplar pool to improve both the task description and few-shot prompt. Then, ask LLM to perform text classification with this prompt.
@@ 3.2 In-process
where $S$ is the set of already selected items that is initially empty, and $\lambda$ is a trade-off parameter between relevance and diversity ($0\leq\lambda\leq 1$). In this way, we retrieve exemplars that are either semantically close to the query text with the same correct label or hard negatives that share some similarity with $\mathbf{x}_{q}$ but from a different class. These input-out pair examples are automatically appended to the task description for further enhancement.
Coarse annotation with weak LLMs.
Using a prompt enhanced with clearer task descriptions and carefully selected examples, we employ chat LLMs to assign labels from predefined options. To efficiently handle the labeling process, we utilize two instances of a weaker LLM (e.g., GPT-3.5) to label all the unlabeled texts twice. This approach mitigates high costs, usage limits, and long processing times. The weaker LLMs are expected to agree on easy queries with high accuracy, while any discrepancies are tracked in a mismatch collection for further review.
3.3 Post-process
In this stage, we leverage more advanced LLMs (e.g., GPT-4) and in-context learning techniques to address mismatches identified in the in-process stage. The prompts used here are refined versions of those from the previous stage, focusing on enhancements for fine-grained annotation and versatility. Although these prompts are less cost-effective, they are applied only to a limited number of queries in the mismatch collection, thereby mitigating the cost concern.
@@ 3.3 Post-process
A chain-of-thought (CoT) prompt guides a large language model (LLM) through a step-by-step reasoning process to enhance its ability to tackle complex tasks (wei2023chainofthoughtpromptingelicitsreasoning). To implement it, the LLM is instructed to first analyze the content according to the task description, providing reasoning at each step before delivering the final answer. This approach works by mimicking human problem-solving, breaking down tasks into smaller components, which helps the model grasp the underlying logic and produce more accurate responses by not only assigning a label but also offering the reasoning behind it. The output of CoT is a sequence $y=(y_{1},y_{2},\dots,y_{t})$ tokens, where $y_{t}$ is the desired prediction and $y_{<t}$ are the reasons:
@@
Advanced LLM as a judge.
Leveraging an LLM as a judge (zheng2023judgingllmasajudgemtbenchchatbot) involves utilizing its advanced reasoning capabilities to evaluate and validate the outputs of other models. In our approach, we first ask two weaker LLMs to provide annotations and reasoning to the small mismatch set through a CoT process. Then, an advanced LLM assesses the quality, accuracy, and consistency of these responses, adding an extra layer of validation. This method refines coarse predictions on challenging examples, enhancing the reliability of outputs in complex tasks.
@@ 3.4 Summary of Framework
In this section, we summarize the entire framework, outlining each module’s input-output and the human involvement required throughout the process.
The “pre-process” stage prepares a pool of representative and diverse examples for human annotation in the subsequent stages. The input is human-collected, unlabeled texts, which we process using LLM embedding models, saving selected indices as an exemplar pool. Following this stage, human experts accurately annotate the texts in the pool and draft an initial prompt that includes only a task description. The “in-process” stage focuses on enhancing this initial prompt with a refined task description and more appropriate examples. Here, two weaker LLMs generate coarse annotations for the unlabeled examples. Users are responsible for running the task description generator, verifying its validity, and appending it to the initial prompt. After labeling, humans clean the predictions, identify mismatches, and record their indices in the dataset. The “post-process” stage refines the predictions for these mismatches and provides reasoning for the adjustments. Humans shall clean the responses from an advanced LLM, replace coarse predictions with fine-grained ones, or conduct human evaluations with the aid of LLM-generated reasoning.
By following this framework, we minimize human effort, enabling rapid experimentation and delivering high-accuracy predictions infused with human knowledge, all while reducing costs in terms of both time and resources.
@@ 4 Experiments
We evaluated our method using three distinct text datasets with human-labeled classifications. In each experiment, we initially treated all texts as unlabeled and tasked LLMs with predicting the labels. We use the model “text-embedding-3-small” from OpenAI to convert all texts into embedding vectors and reduce the size of a vector to 24 using UMap. Subsequently, we selected 80 exemplars through KMeans from each dataset to form the example pool used to develop the enhanced prompts [^3] We conducted hyperparameter tuning experiments to assess the impact of varying the number of exemplars, testing sample sizes of 20, 40, 60, 80, and 100. Although F1 scores consistently improved as the number of exemplars increased, the rate of improvement slowed beyond a certain point. For our analysis, we chose to use only 80 exemplars in all three experiments, deliberately demonstrating that our method can outperform traditional ML approaches even with a relatively modest number of examples. As a result, the performance metrics we report are conservative estimates, not reflecting the highest potential performance. Full results from the hyperparameter experiments can be found in Appendix 1. Throughout our experiments, “GPT-3.5-turbo” and “Mistral-medium” were consistently employed as the weak models. A $\lambda$ parameter of value 0.8 is chosen for the MMR dynamic retriever. When referencing the advanced or judge model, we refer to either “GPT-4-turbo” or “Mistral-large-latest”.
In the first experiment, using a clean, correctly labeled multi-class news dataset(greene06icml), we demonstrated our method’s exceptional performance in accurately labeling text topics, even for lengthy inputs. Our three-stage approach systematically enhanced precision, with each module contributing cumulative improvements, as confirmed by an ablation study. Notably, our method achieved a significant improvement in overall F1 score, increasing from 0.94 with the naïve LLM labeling to an impressive 0.96.
In the second experiment, we specifically highlight the effectiveness of our second-stage generator. We tasked LLMs with labeling tweets as ”support” or ”oppose” in relation to Brett Kavanaugh’s Supreme Court confirmation process. Initially, we observed that weaker LLMs often struggled with discerning sentiment and stance, leading to confusion. However, when utilizing our second-stage in-process generator, the LLM generated an enhanced prompt by summarizing the labeling rules for ”support” and ”oppose” based on 80 human-labeled tweets. This resulted in a dramatic improvement in labeling accuracy, with F1 scores soaring from 0.57 to 0.95.
In the third experiment, we attempted to replicate Fowler et al.’s (2021) study, which examines how the medium of campaign ads influences their tone. However, we encountered a very noisy, human-labeled dataset. Our method effectively uncovered potential issues within this supposedly gold-standard dataset, highlighting how noisy human labels can significantly impact downstream political science analysis. We also addressed some pitfalls of using our method, emphasizing the importance of careful data handling.
4.1 Classifying BBC News Reports Topics
In our first experiment, we applied our method to a multi-category classification task involving extensive and lengthy text. We chose to label the topics of BBC news reports due to their diversity and relevance in benchmarking machine learning models. The dataset comprises 2,225 news articles sourced from the BBC News website, covering stories across five topical areas—business, entertainment, politics, sport, and tech—from the years 2004-2005. This dataset, originally compiled by Greene and Cunningham (2006), has been widely used in machine learning research as a benchmark for evaluating the performance of various classification algorithms. The diversity and structure of this dataset make it an ideal candidate for testing the robustness and accuracy of our proposed method in a real-world, multi-class classification scenario.
Starting with a simple heuristic prompt, as depicted in Figure 2, we guided two weak LLMs to identify the primary topic of a news report from five categories: politics, business, sport, entertainment, and technology. Next, we employed a prompt generator, feeding the LLM with 80 exemplars labeled with accurate human classifications and asking the LLM to summarize the rules for categorizing news reports. This newly generated prompt was then fed back to the weak LLMs. As a result, the models not only corrected some of their initial misclassifications but also provided justifications for their decisions. In both zero-shot and five-shot settings [^4] We employed five-shot prompting, providing the LLMs with one example per category since there are five topics to classify., these weaker models showed significant improvement in performance when we dynamically selected 80 exemplars compared to their initial raw predictions [^5] Detailed results are provided in Tables LABEL:table:Topic_0shot and LABEL:table:Topic_5shot.. For example, as illustrated in Figure 3 when using zero-shot prompting, both GPT-3.5 and Mistral-medium initially achieved F1 scores of approximately 0.89 in the ”politics” category. However, these scores increased to around 0.94 after an enhanced prompt was generated. In categories like ”sport,” where the LLMs already performed exceptionally well with the naïve prompt—achieving F1 scores above 0.97—our method further boosted their performance. we were able to raise the F1 scores to around 0.99 putting into the enhanced prompt.
In Figure 2, we present an example where the two weak LLMs continued to differ in their classifications even after prompt enhancement [^6] The full text of this news report example is available in the Appendix LABEL:fig:Yahoo_News_Report. In this case, one model classified the news report as related to technology, while the other identified it as business. To resolve this discrepancy, we applied our third stage, which involved chain-of-thought prompting combined with a judge model. “GPT-4-turbo”, as the judge model, ultimately classified the report as business, aligning with the human label. This decision was well-justified, as the judge model highlighted that the report, while contextualized within technology, primarily focused on business aspects.
Figure 2: An Example of 0-Shot Prompts for Analyzing BBC News Reports
In your judgment, what is the major topic of the given news report Text? Please choose only one answer from the following options: ”politics”, ”business”, ”sport”, ”technology”, or ”entertainment”. Complete the task very succinctly using only one word written between ’<’ and ’>’. [Enhanced Prompt][BBC News Text (lengthy text in the Appendix LABEL:fig:Yahoo_News_Report)] Weak Chat Model1: The news report discusses the history and future of Yahoo, including its competition with Google and MSN, its acquisitions, and its plans for the future. It also mentions the importance of search technology and the potential expansion into rich media. Based on this, the major topic of the news report is ”technology.” Therefore, the correct answer is ¡technology¿. Weak Chat Model2: The text discusses the future prospects of Yahoo, a major internet company, focusing on its competition with Google and MSN, its business strategies, and its potential dominance in the next decade. Therefore, the correct answer is ¡business¿. Judge Prompt:

### v2

@@ 3.1 Data Preprocessing
LLMs convert natural language into fixed-size embedding vectors, enabling comparison of semantic similarity. We apply this transformation to all unlabeled texts using models such as OpenAI’s ”text-embedding-3-small.” Following (Steck_2024), we use cosine distance to quantify similarity between embeddings:
@@
Uniform Manifold Approximation and Projection (UMAP) is a widely used dimensionality reduction technique that preserves relationships in high-dimensional data. Conceptually, it builds on traditions in political science like factor analysis (kim1978introduction), multidimensional scaling (borg2007modern), and correspondence analysis (greenacre2017correspondence), which have been used to uncover latent dimensions in political behavior, such as ideological scaling from roll-call votes (poole1985spatial) and survey data (aldrich1977method). UMAP advances these techniques using non-linear methods, enabling detection of complex structure in text embeddings while remaining compatible with established scaling approaches.
In practice, UMAP is also widely supported in computer science, especially for high-dimensional data. We use UMAP (mcinnes2020umapuniformmanifoldapproximation) with cosine distance to reduce text embeddings, improving efficiency in later stages. The reduced representations also enable Euclidean distance calculations, which are faster and more broadly supported in downstream applications. For further implementation details, see Appendix C. [^5] While we use UMAP here, our method supports alternative techniques such as Principal Component Analysis (PCA) (wold1987principal), t-SNE (maaten2008visualizing), and Independent Component Analysis (ICA) (hyvarinen2000independent), depending on dataset and task requirements.
3.2 Automatic Prompt Optimization
After structuring the dataset through embedding generation and feature reduction, we apply automatic prompt optimization. LLMs face a significant constraint in their context window size, which limits the number of tokens  [^6] In a Large Language Model (LLM), a ”token” is the smallest unit of text that the model processes, essentially representing a word or a sub-word, which is used as the building block for understanding and generating language the model can process simultaneously in few-shot learning scenarios. Moreover, within this window, earlier tokens in a long sequence may receive less attention, further constraining the effective use of examples for inference. This limitation necessitates a method that not only selects exemplars that are highly relevant and representative of the broader dataset but also minimizes the number of tokens used. Such a method can mitigate the risk of earlier tokens being underweighted, thereby preserving their contribution to inference. To achieve this, we propose a novel approach: first, we select a pool of $M$ representative examples for human labeling, where $k<M\ll N$, with $N$ representing the total number of examples under study and $k$ denoting the number of exemplars included in the context window for few-shot in-context learning. Then, at inference time, when calling the commercial API to classify unlabeled query text, we dynamically select the most relevant $k$ examples to augment the prompt. In the automatic prompt optimization stage, we systematically curate this exemplar pool for human labeling. This curated set serves two critical purposes: first, it provides the LLM with a structured reference to infer the underlying labeling rules, reducing ambiguity in classification. Second, it enables an adaptive selection mechanism, ensuring that the most contextually relevant examples are incorporated for each query, thereby enhancing classification accuracy and robustness. By creating this initial pool of human-annotated examples, we lay the foundation for more effective and efficient use of the LLM in subsequent stages, balancing the need for comprehensive knowledge injection with the constraints of the model’s context window. In this stage, we prepare this exemplar pool, and here are the steps.
Labeling Sample Selection.
[…]
@@ 3.2 Automatic Prompt Optimization
Figure 1: Exemplar Selection Process for Few-Shot Learning: Demonstrating k-means Clustering to Identify Representative Samples for Labeling
Our automatic prompt optimization method moves beyond prompt engineering by creating refined prompts based on human-labeled exemplars, rather than relying on intuition or trial and error. [^8] Figure A2 in the Appendix illustrates two workflows for finding optimal prompts in LLM applications: Prompt Engineering and our Automatic Prompt Optimization method. A detailed comparison between the two approaches is also available in Table A3 in the Appendix, highlighting their respective processes. Prompt engineering starts with an initial prompt, which is then tested in LLMs to generate results. These results are manually checked by researchers for consistency and accuracy. If accuracy is low, the researcher enters a ”Revise and Try” loop, making adjustments based on intuition. This trial-and-error process is time-consuming and lacks clear feedback on where and how the prompt may have fallen short. As a result, creating a final satisfactory prompt often requires multiple speculative revisions, making it inefficient and hard to replicate.
In contrast, our automatic prompt optimization method offers a systematic and transparent approach by using human-labeled exemplars rather than relying on intuition or trial and error. This process begins with researchers labeling a small set of curated examples, typically fewer than 100, which serve as a foundation for the model to generate an enhanced prompt tailored to the specific task criteria. This optimized prompt undergoes human verification, where it is assessed not only by numerical performance metrics, such as F1 scores, but also by its ability to accurately reflect human-labeled criteria and consistently produce reliable outputs. By grounding the prompt on well-defined exemplars, this workflow eliminates guesswork and minimizes trial and error. This structured, exemplar-based approach ensures effective and transparent prompt refinement.
3.3 Dynamic Exemplar Selection
In this stage, to address the limitations of using fixed exemplars in few-shot learning and their potential negative impact on classification outcomes, we propose a method to ensure that exemplars are both relevant to the specific query text and representative of the broader dataset. This dynamic selection process enhances the effectiveness of few-shot learning by tailoring exemplars to the specific input while maintaining comprehensive coverage of the dataset.
@@ 3.3 Dynamic Exemplar Selection
where $S$ is the set of already selected items that is initially empty, and $\lambda$ is a trade-off parameter between relevance and diversity ($0\leq\lambda\leq 1$). In this way, we retrieve exemplars that are either semantically close to the query text with the same correct label or hard negatives that share some similarity with $\mathbf{x}_{q}$ but from a different class. 1 illustrates a scenario where a query text is to be labeled, and the MMR (Maximal Marginal Relevance) algorithm selects two representative exemplars from the pool of five previously selected exemplars. This approach, known as two-shot learning, identifies the most relevant and similar exemplars to the query text, ensuring a contextually informed labeling process for few-shot learning.
Coarse Annotation with a Consensus Mechanism.
Using a prompt enhanced with clearer task descriptions and carefully selected exemplars, we employ two LLMs to assign labels from predefined options. To improve accuracy and ensure reliability, we run the labeling process twice using two separate LLMs, allowing for comparison and verification of results. This approach reduces labeling inconsistencies by capturing high-confidence agreements while identifying hard cases where the models disagree. Discrepancies are tracked in a mismatch collection for further review, making the results more robust and interpretable.
3.4 Consensus Mechanism
In this stage, we utilize LLMs with in-context learning techniques to address mismatches identified in the dynamic exemplar selection stage. The prompts used here are refined versions of those from the previous stage, incorporating enhancements for fine-grained annotation and versatility. While these prompts may be computationally intensive, they are applied only to a limited number of queries in the mismatch collection, ensuring cost efficiency.
@@ 3.4 Consensus Mechanism
A chain-of-thought (CoT) prompt guides a large language model (LLM) through a step-by-step reasoning process to enhance its ability to tackle complex tasks (wei2023chainofthoughtpromptingelicitsreasoning). To implement it, the LLM is instructed to first analyze the content according to the task description, providing reasoning at each step before delivering the final answer. This approach works by mimicking human problem-solving, breaking down tasks into smaller components, which helps the model grasp the underlying logic and produce more accurate responses by not only assigning a label but also offering the reasoning behind it. CoT prompting has emerged as one of the most successful methods for improving the reasoning and inferential abilities of LLMs, making it a cornerstone of advanced prompt engineering (kojima2022large). The output of CoT is a sequence $y=(y_{1},y_{2},\dots,y_{t})$ tokens, where $y_{t}$ is the desired prediction and $y_{<t}$ are the reasons:
@@
LLM-Based Consensus Arbitration.
An LLM can serve as a judge (zheng2023judgingllmasajudgemtbenchchatbot), leveraging its reasoning capabilities to evaluate and validate the outputs of other models. In our approach, two LLMs first annotate and reason through a small mismatch set using a Chain-of-Thought (CoT) process. The judging LLM then analyzes the reasoning steps provided by both models, assessing the quality, accuracy, and consistency of their responses. By evaluating the justification behind each prediction, rather than just the final label, this process enhances interpretability and ensures a more rigorous validation. This arbitration mechanism mirrors a multi-coder scenario, where two coders may disagree, and a third adjudicates discrepancies to ensure classification reliability. Previous research has demonstrated that incorporating LLMs as evaluators improves performance in high-precision tasks requiring structured reasoning and interpretability (zheng2023judgingllmasajudgemtbenchchatbot). By integrating CoT reasoning with systematic judgment, our approach enhances consistency, reduces ambiguity, and strengthens the robustness of model outputs.
@@ 3.5 Summary of Framework
In this section, we summarize the entire framework, outlining each module’s input-output and the human involvement required throughout the process as illustrated in Figure 2. The “automatic prompt optimization” stage prepares a pool of representative and diverse examples for human annotation in the subsequent stages. The input is human-collected, unlabeled texts, which we process using LLM embedding models, saving selected indices as an exemplar pool. Following this stage, human experts accurately annotate the texts in the pool and draft an initial prompt that includes only a task description. The “dynamic exemplar selection” stage focuses on enhancing this initial prompt with a refined task description and more appropriate examples. Here, two LLMs generate coarse annotations for the unlabeled examples. Users are responsible for running the task description generator, verifying its validity, and appending it to the initial prompt. After labeling, humans clean the predictions, identify mismatches, and record their indices in the dataset. The consensus mechanism then refines the predictions for these mismatches and provides reasoning for the adjustments. Humans review and clean the responses, replacing coarse predictions with fine-grained labels or conducting manual evaluations supported by LLM-generated reasoning.
By following this framework, we first leverage the in-context learning capabilities of large language models (LLMs), eliminating the need for intensive feature engineering, large labeled training datasets, task-specific training or fine-tuning, and high computational resources like GPUs. Therefore, this method addresses the limitations of commonly used text classification methods in political science. Second, we overcome two key limitations of LLMs: their heavy reliance on well-crafted prompts and the static nature of exemplars in few-shot learning. Our method dynamically optimizes prompts and selects relevant exemplars, making LLMs a more powerful and reliable tool for text classification in political science research.
Figure 2: Overview of the three core stages of our PoliPrompt framework
@@ 4 Experiments
Overview.
We evaluated our method across three classification tasks in political science: sentiment analysis and stance detection on Twitter data related to Brett Kavanaugh’s Supreme Court confirmation, and multi-category tone classification for campaign ads from the 2018 U.S. elections. The Twitter experiments demonstrate that our method can reliably distinguish between closely related concepts such as sentiment and stance—an area where many classifiers traditionally struggle. Furthermore, the campaign ad experiments highlight how the consensus mechanism enhances label consistency and helps detect potential noise in human annotations, particularly in more subjective or complex multi-category settings.
For sentiment and stance detection, Twitter was chosen for several key reasons. First, as one of the most widely used social media platforms, Twitter provides a wealth of real-time information during major events. For example, it has been a primary source of breaking news, especially during recent U.S. Presidential Elections (allcott2017social). The unstructured data on Twitter offers valuable insights into public opinion and attitudes toward political figures and events (evans2014twitter). Second, sentiment analysis and stance detection are crucial in political science for understanding public views on significant political topics or figures (hopkins2010method, grimmer2013text, gentzkow2019measuring). Studies like those by hopkins2010method and bond2015quantifying show how analyzing sentiment and stance can reveal shifts in political attitudes and patterns of polarization. Finally, applying sentiment analysis and stance detection to the same dataset presents unique challenges, especially in distinguishing between the two. bestvater2023sentiment argue that classifiers trained for sentiment analysis often struggle with stance classification, highlighting the challenge of distinguishing these frequently conflated political concepts. Our goal is to test whether our method can overcome this challenge, effectively separating sentiment from stance with high reliability. Specifically, we aim to achieve accurate stance detection through prompt optimization alone, without the need for retraining.
We further applied our method to a multi-category classification task- analyzing campaign ads from the 2018 elections, categorizing them as contrast, attack, or promote. This task was chosen for several reasons. First, substantively, analyzing campaign ads provides valuable insights into campaign strategies and behavior across different media, including social media and traditional broadcast (fowler2021political). Second, this task allows us to evaluate our method’s performance in multi-class classification, expanding beyond binary classification. Third, the longer text format of campaign ads compared to social media posts adds a layer of complexity, testing our method’s capacity to handle extended text [^9] We conducted one last experiment involving multi-category topic classification on lengthy BBC news reports. This experiment demonstrates that our method is versatile and effective for classifying various types and lengths of text data, extending beyond short texts such as tweets. Detailed methodology and results are available in  E. .
@@
Each dataset used in our experiments was manually labeled, serving as the ground-truth benchmark. In our experiments, we initially treated all texts as unlabeled, tasking the LLMs with predicting labels using our proposed method. We then compared the predicted labels to the human-labeled benchmark to evaluate the accuracy of our approach. For each dataset, we used the ”text-embedding-3-small” model from OpenAI to convert all texts into embedding vectors. To reduce the dimensionality and improve efficiency, we applied UMAP to shrink each vector to 24 dimensions. We then selected 80 exemplars from each dataset using k-Means clustering. These exemplars formed the example pool, which was leveraged to generate optimized prompts for LLM-based text classification [^10] We conducted hyperparameter tuning experiments to assess the impact of varying the number of exemplars, testing sample sizes of 20, 40, 60, 80, and 100. Although F1 scores consistently improved as the number of exemplars increased, the rate of improvement slowed beyond a certain point. For our analysis, we chose to use only 80 exemplars in all experiments, deliberately demonstrating that our method can outperform traditional ML approaches even with a relatively modest number of examples. As a result, the performance metrics we report are conservative estimates, not reflecting the highest potential performance. Full results from the hyperparameter experiments can be found in Appendix 3.. In our analysis, we define unoptimized application as the direct application of LLMs without our optimization techniques and refer to their corresponding output as unoptimized results. In contrast, results produced through the first two stages of our method—incorporating automatic prompt optimization and dynamic exemplar selection—are referred to as optimized results for Mistral and GPT models.
In our experiments on sentiment analysis and stance detection, we applied our method to classify over 3,660 tweets related to Brett Kavanaugh’s Supreme Court confirmation, labeling each tweet as either ”positive” or ”negative” for sentiment and ”support” or ”oppose” for stance. We chose this task because prior research highlights the challenges of distinguishing sentiment from stance, as emphasized in the seminal article ”Sentiment is Not Stance: Target-Aware Opinion Classification for Political Text Analysis” (bestvater2023sentiment). By using automatic prompt optimization and supplying the model with 80 exemplars labeled for both sentiment and stance, we demonstrated that LLMs can effectively generate and refine clear definitions for these two dimensions. Additionally, the optimized prompt includes illustrative examples where sentiment and stance diverge, providing researchers with a reliable reference for crafting prompts that classify sentiment and stance independently. For researchers aware of the confusion surrounding concepts like sentiment and stance, our method shows that LLMs can effectively differentiate and summarize classification rules for these concepts when provided with a limited set of labeled exemplars. This demonstrates a significant advantage in transparency and explainability for researchers.
For sentiment classification, we implemented our three-stage method. Initially, the unoptimized LLMs, yielded low F1 scores. After applying automatic prompt optimization and dynamic exemplar selection, we observed a marked improvement, with F1 scores increasing by over 15%. In the final stage, our consensus mechanism further enhanced accuracy, pushing F1 scores to nearly 92%. These findings show that directly applying LLMs for political text classification often leads to suboptimal results, while each stage of our method delivers measurable gains in accuracy. Most importantly, the transparency offered by prompt optimization allows researchers to pinpoint possible sources of error in the prompts and revise them precisely, without relying on trial-and-error associated prompt engineering.
Following sentiment classification, we applied our method to stance detection, highlighting our approach’s effectiveness in classifying complex, often conflated concepts. Initially, the LLMs struggled to differentiate sentiment from stance, resulting in frequent misclassifications. Through automatic prompt optimization, however, the models dynamically generated an enhanced prompt that clarified the distinctions between ”support” and ”oppose,” based on a set of 80 human-labeled tweets. This refinement yielded a dramatic improvement, with F1 scores increasing from 57% to 95%.
In our final experiment, we attempted to replicate Fowler et al.’s (2021) study on how the medium of campaign ads influences tone. However, we encountered a highly noisy, human-labeled dataset, where labeling inconsistencies among human annotators became apparent. Our third stage of consensus mechanism effectively identified these conflicts, particularly in cases involving challenging or conceptually ambiguous classifications. This experiment highlights the significant impact a noisy dataset can have on downstream political analysis and demonstrates how our method can serve as a safeguard, flagging problematic labels and helping to prevent unreliable estimates early in the analysis process.
4.1 Sentiment Analysis and Stance Detection toward Brett Kavanaugh’s SCOTUS Nomination Using Twitter Data
In their 2023 study, Bestvater and Monroe analyzed a hand-coded dataset of 3,660 tweets, labeling each for both sentiment (positive or negative emotional tone) and stance (support or oppose regarding an issue). To evaluate the ability of large language models (LLMs) to differentiate these two closely related dimensions, we selected a subset of 80 tweets from this dataset, containing both sentiment and stance labels, as input for automatic prompt optimization. This setup tested whether the LLM, guided by our approach, could generate an optimized prompt that effectively clarifies sentiment and stance.
Our initial prompt was straightward and unoptimized, as illustrated in Figure 3. Through automatic prompt optimization, our method transformed this prompt into an enhanced version with guidelines, which not only provides definitions of sentiment and stance but also provides four clear examples illustrating different scenarios. This approach goes beyond basic instructions by integrating guidelines on special cases that commonly lead to confusion between sentiment and stance, such as instances where criticism of political opponents (rather than the primary subject, Brett Kavanaugh) could imply a negative sentiment without necessarily indicating opposition to Kavanaugh’s confirmation. By providing specific instructions for cases where sentiment and stance may diverge, this prompt allows the LLM to handle complex examples, such as sarcasm or irony, more effectively.
This prompt provides clear, interpretable distinctions between stance and sentiment, offering a reliable reference for researchers who may want to classify these concepts independently in future work. The prompt allows researchers to easily verify clarity in definitions, reducing the need for trial-and-error adjustments and enabling consistent, accurate classification.
Figure 3: Comparison of Simple Heuristic and Enhanced Prompts for Sentiment and Stance Classification.
Comparison of Simple Heuristic and Enhanced Prompts Simple Heuristic Prompt Determine the sentiment and stance of the Twitter text related to Brett Kavanaugh’s Supreme Court confirmation. Ensure your evaluation clearly differentiates between the emotional sentiment and the political stance.   Enhanced Prompt You are given Twitter text about Brett Kavanaugh’s confirmation. Your task is to determine both the sentiment (positive or negative) and the stance (support or oppose). Guidelines: Sentiment: Assess the emotional tone, either positive (e.g., approval) or negative (e.g., condemnation), based on the language used. Stance: Identify the political position regarding Kavanaugh’s confirmation, either support (e.g., advocating for or defending the confirmation) or oppose (e.g., arguing against or resisting the confirmation). Note: Sentiment and stance may not always align. For example, a positive emotional tone may still reflect an opposing stance, while a negative tone could support confirmation. Key Indicators: Positive Sentiment: Words indicating happiness or affirmation. Negative Sentiment: Words indicating condemnation or anger. Support Stance: Text that defends, advocates for, or agrees with Kavanaugh’s confirmation. Oppose Stance: Text that criticizes, argues against, or resists Kavanaugh’s confirmation. Special Cases: Humor or lighthearted criticism of the confirmation process itself, rather than the nominee, may indicate a positive sentiment but an opposing stance. Criticism of Kavanaugh’s opponents or the process, rather than Kavanaugh himself, may suggest negative sentiment but a supportive stance. Examples: Positive, Support: ”Kavanaugh is a great choice, he should be confirmed.” Positive, Oppose: ”Glad to see the protests against Kavanaugh.” Negative, Support: ”This investigation is unfair, but Kavanaugh will be confirmed.” Negative, Oppose: ”Kavanaugh’s actions are disqualifying, he should not be confirmed.”
Note: Red sections represent definitions and explanations for sentiment, while blue sections represent stance. This figure illustrates how, when prompted to treat sentiment and stance as distinct concepts, LLMs—guided by 80 carefully selected labeled exemplars—successfully differentiate between these two dimensions. The LLM-generated prompt provides detailed guidelines and examples, demonstrating a clear understanding of the unique characteristics of sentiment and stance.
In the next step, we used the same dataset as our ground-truth benchmark and tasked both GPT-3.5-turbo and Mistral-medium-latest with classifying only sentiment, starting with an unoptimized heuristic prompt based on the original instructions given to human coders by Bestvater and Monroe (bestvater2023sentiment). This unoptimized prompt is illustrated in Figure 4.
Figure 4: Prompts for Sentiment Analysis about Brett Kavanaugh
Prompts for Sentiment Analysis towards Brett Kavanaugh Unoptimized Heuristic Prompt:
In your judgment, whether the specific sentiment is positive or negative? Please choose your answer only from the 2 options – ”positive” and ”negative”. Complete the task very succinctly using only one word written between ’<’ and ’>’.
  Enhanced Prompt:
In your judgment, whether the specific sentiment is positive or negative? Please choose your answer only from the 2 options – ”positive” and ”negative”. Complete the task very succinctly using only one word written between ’<’ and ’>’. Evaluate the overall emotional tone, context, and specific language to determine if the sentiment is positive or negative.
negative:
Choose ¡negative¿ when the text expresses opposition condemnation, criticism, dissatisfaction, frustration, anger, or uses negative language such as expletives, derogatory terms, or focuses on conflict, harm, or unethical behavior.
positive:
Choose ¡positive¿ when the text shows approval affirmation, support, or uses favorable adjectives.
Note: Italicized sections were generated after applying automatic prompt optimization. The enhanced prompt succinctly summarizes the criteria for labeling text as ”positive” or ”negative,” providing clarity and reducing ambiguity in sentiment classification.
Using this unoptimized prompt, GPT-3.5-turbo and Mistral-medium-latest achieved moderate F1 scores of 84% and 88%, respectively, as illustrated in Figure 5. We then applied automatic prompt optimization. This enhanced prompt in Figure 4 provided clearer criteria for identifying sentiment, focusing on emotional tone to improve classification precision. The optimized prompt increased GPT-3.5-turbo’s F1 score by over 5% and Mistral’s by 3%. When both models aligned in their predictions, the F1 score reached 94%, as shown in the Agreement columns in Figure 5. A third-stage consensus mechanism further refined classification reliability, highlighting the importance of structured prompt optimization and exemplar selection. This process illustrates the limitations of using unoptimized LLM outputs.
A notable outcome in applying our automatic prompt optimization method is the transparency it provides in distinguishing between stance and sentiment in prompts. Initially, the enhanced prompt generated for sentiment analysis included terms like “text expresses opposition, criticism…” for “negative” and “text shows approval, support, or uses favorable language” for “positive.” Even after using 80 representative, human-labeled exemplars, the LLMs continued to confuse stance with sentiment.
Thanks to the transparency of our method, we were able to directly examine the enhanced prompt, pinpoint the source of confusion, if any, and adjust the language accordingly. Specifically, we revised terms from “opposition” to “condemnation” and “approval” to “affirmation,” guiding the model to focus on emotional tone rather than political stance. This refinement aligned the prompt with the definitions established in our previous application of automatic prompt optimization, as shown in Figure 3, where positive sentiment encompasses happiness and affirmation, and negative sentiment includes terms indicating condemnation and anger. This example of sentiment classification highlights a clear advantage of our approach: by generating prompts grounded in structured, task-specific logic, our method allows researchers to systematically inspect and validate prompts without relying on the blind trial-and-error cycles typical of traditional prompt engineering. This transparency ensures that sentiment classification remains centered on emotional tone, enhancing interpretability, accuracy, and explainability by allowing researchers to understand what qualifies texts as positive or negative, including the specific emotions they express and the reasoning behind each classification.